# Plan2Align: Predictive Planning Based Test-Time Preference Alignment in Paragraph-Level Machine Translation

### v1

@@ Abstract
Machine Translation (MT) has been predominantly designed for sentence-level translation using transformer-based architectures. While next-token prediction based Large Language Models (LLMs) demonstrate strong capabilities in long-text translation, non-extensive language models often suffer from omissions and semantic inconsistencies when processing paragraphs. Existing preference alignment methods improve sentence-level translation but fail to ensure coherence over extended contexts due to the myopic nature of next-token generation. We introduce Plan2Align, a test-time alignment framework that treats translation as a predictive planning problem, adapting Model Predictive Control to iteratively refine translation outputs. Experiments on WMT24 Discourse-Level Literary Translation show that Plan2Align significantly improves paragraph-level translation, achieving performance surpassing or on par with the existing training-time and test-time alignment methods on LLaMA-3.1 8B.
Plan2Align: Predictive Planning Based Test-Time Preference Alignment
in Paragraph-Level Machine Translation
Kuang-Da Wang1,† Teng-Ruei Chen1,† Yu Heng Hung1 Shuoyang Ding2 Yueh-Hua Wu2 Yu-Chiang Frank Wang2Chao-Han Huck Yang2Wen-Chih Peng1Ping-Chun Hsieh1 1National Yang Ming Chiao Tung University  2NVIDIA {gdwang.cs10@,raychen.cs12@,wcpeng@cs.,pinghsieh@}nycu.edu.tw{hucky,krisw,shuoyangd}@nvidia.com equal contribution
@@ 1 Introduction
Machine Translation (MT) has primarily leveraged transformer-based encoder-decoder neural architectures, as exemplified by models like M2M-100 (Fan et al., 2021) and MT5 (Xue et al., 2021). However, the emergence of Large Language Models (LLMs), such as the GPT series (Brown et al., 2020; Achiam et al., 2023), LLaMAs (Touvron et al., 2023a, b), and Gemma (Team et al., 2024), has demonstrated remarkable efficacy in a wide range of NLP tasks (Stiennon et al., 2020; Hendrycks et al., 2021; Srivastava et al., 2023; Yu et al., 2024; Zhong et al., 2024), including multilingual translation. This progress has led to increasing interest in leveraging LLMs for MT, particularly for handling long-form text. LLMs excel at long-text translation due to their ability to model dependencies across extended contexts, enabling improved coherence and fluency. Models such as Tower (Alves et al., 2024) and GPT-4o (OpenAI, 2024) have demonstrated state-of-the-art performance, surpassing traditional MT systems in both fluency and contextual consistency.
Figure 1: Machine translation via predictive planning from the perspective of trajectory optimization.
However, non-extensive (i.e., smaller than 10B) LLMs often struggle with long-text translation, exhibiting sentence omissions, semantic inconsistencies, and hallucinations (Wu et al., 2024). These limitations stem from the myopic nature of next-token prediction, where local fluency is prioritized at the cost of global coherence, leading to degraded translation quality in extended discourse. One-pass translation remains inherently limited, as it lacks iterative refinement, making it prone to over-translation, under-translation, and compounding errors (Mandya et al., 2020). While traditional MT systems incorporate techniques such as re-ranking, iterative decoding, and constrained beam search to improve coherence and fidelity, LLM-based MT operates purely in a next-token prediction paradigm, lacking an explicit self-correction mechanism (Hadj Ameur et al., 2019; Feng et al., 2024). As a result, errors introduced early in the generation process propagate across extended text spans, leading to semantic drift, loss of contextual dependencies, and degraded translation quality in long-form content.
Recent research has investigated preference alignment as a means to improve translation quality in small-scale LLMs, leveraging preference data to guide optimization. Methods such as CPO (Xu et al., 2024b) and WAP (Wu et al., 2024) have shown promising results in aligning translations to preferences. However, existing methods face two fundamental challenges in long-text translation:
@@
    Preference alignment techniques operate at specific granularity levels; when applied at the long-text level, they often struggle to preserve sentence coherence across paragraphs.
@@
    Unlike training-time alignment, which requires costly fine-tuning, testing-time alignment methods must incorporate a technique to recover from past errors or low-quality answers caused by compounding errors during inference. This is crucial because these past mistakes can lead to further errors, as earlier responses influence the generation of answers to subsequent questions (Mandya et al., 2020).
To address these challenges, we propose Plan2Align, a test-time alignment method that employs a self-rewriting framework based on the principle of Model Predictive Control (MPC), a classic predictive planning method for trajectory optimization (Rao, 2014) widely used in robotics (Tassa et al., 2012, 2014; Nagabandi et al., 2020) and motion planning (Ji et al., 2016). Rather than fine-tuning LLMs, Plan2Align iteratively refines translations by integrating high-quality segments from previous iterations at test time, enhancing coherence and fluency in long-text translation.
This study focuses on paragraph-level MT, where translation quality is evaluated beyond isolated sentences. Specifically, we define context-level as a group of adjacent, semantically related sentences, while a paragraph consists of multiple interconnected contexts. Ensuring high-quality translation at the context level is crucial for maintaining coherence and accuracy at the paragraph level. Specifically, this study is twofold: (i) to develop a test-time alignment method that matches or surpasses fine-tuning approaches, and (ii) to enhance paragraph-level translation by improving consistency and fidelity across context-level translations.
As a predictive planning method, MPC optimizes a sequence of decisions by continuously adjusting future actions based on past observations (Camacho and Bordons, 2007; Kouvaritakis and Cannon, 2016). Inspired by this, we draw an analogy between trajectory optimization and translation optimization, where prior translations inform subsequent refinements, and recast MT as an iterative optimization process rather than a one-pass mapping. Figure 1 illustrates this correspondence, highlighting the intuition behind our approach. While LLMs excel in contextual understanding, their one-pass translations remain suboptimal. By iteratively leveraging high-quality past translations, our method enhances coherence and discourse integrity, improving overall quality.
However, directly applying MPC to paragraph-level MT is impractical when entire paragraph-level translation histories are fed into small-scale LLMs due to their limited contextual capacity. To adapt MPC to paragraph-level MT, we introduce two novel mechanisms: (i) Model-predictive context alignment: We modify MPC to selectively retain high-quality contexts from multiple paragraph-level translations with the help of a context buffer, instead of relying on a single best trajectory. This ensures effective accumulation and utilization of translation experience for aligning source and translation contexts. (ii) Self-rewriting tasks: We redefine MT as a self-rewriting task, a structured prompting mechanism that enhances discourse coherence, enables each iteration to build upon prior high-quality translations, and ultimately improves fluency and consistency. By integrating these two strategies, Plan2Align achieves efficient translation refinement within a limited number of iterations.
To validate our approach, we apply Plan2Align to LLaMA-3.1 8B on the WMT’24 Discourse-Level Literary Translation Benchmark, using 6K paragraph-level preference data for test-time alignment. Our method demonstrates substantial improvements, achieving performance on par with or surpassing training-time alignment methods while outperforming existing test-time alignment approaches. Plan2Align bridges the gap in LLM-based paragraph-level translation, introducing a novel test-time alignment framework that ensures high translation quality while maintaining context-level consistency. Moreover, our test-time alignment method can augment other fine-tuning approaches, leveraging a small number of iterations at test time to achieve even better translation performance.
@@ 2 Related Work
2.1 Document-Level Machine Translation
Document-level neural machine translation can be broadly divided into two categories: the sentence-to-sentence (sen2sen) approach and the document-to-document (doc2doc) approach (Maruf et al., 2022; Wang et al., 2023). Translation on a sentence-by-sentence basis, even with augmented contexts, remains inadequate for fully capturing coherence and consistency within texts (Fernandes et al., 2023). Various methods demonstrate the importance of context in translation; (Wu et al., 2023) process context within different components of the same encoder, while (Voita et al., 2018) use attention distributions to integrate context awareness. We are exploring a translation method that effectively bridges the gap between sentence-level and document-level approaches, aiming to harness more contextual information than sentence-level translation while avoiding the extensive resource demands of full document-level translation.
2.2 Preference Alignmnet for LLMs
Large language models (LLMs) often misalign with human preferences. Traditional approaches such as supervised fine-tuning (SFT) (Ziegler et al., 2019), reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), direct policy optimization (DPO) (Rafailov et al., 2023), and simple preference optimization (SimPO) (Meng et al., 2024) refine models post-training but demand significant computational resources. Test-time alignment offers a lightweight alternative by modifying outputs during inference without altering model parameters. Prior methods include InferAligner (Wang et al., 2024b) for cross-model guidance, ARGS (Khanov et al., 2024) with dynamic reward-based adjustment, TreeBoN (Qiu et al., 2024) leveraging tree-search in Best-of-N Sampling, and RAIN (Li et al., 2024) and URIAL (Lin et al., 2024) employing advanced prompt engineering. However, these approaches struggle when early responses are suboptimal, hindering subsequent outputs. To address this, Plan2Align introduces an MPC framework that enables predictive planning, effectively accumulating and leveraging high-quality experience to optimize response generation. By integrating structured foresight into test-time alignment, Plan2Align mitigates early-stage errors, leading to more reliable and adaptable model outputs.
Figure 2: The framework of Plan2Align. To leverage past high-quality translations for preference alignment, Plan2Align transforms the translation task into a self-rewriting process. The LLM refines its translations by revisiting previous outputs, segmenting the paragraph-level result, and aligning at the context level. A reward model, trained on preference data, selects optimal contexts to update the context buffer. Through iterative refinement, the better context buffer improves context quality and generates better rewriting inputs, achieving both paragraph-level preference alignment and context-level coherence.
@@ 3 Methodology
In this section, we introduce the Plan2Align framework (Figure 2). We describe how predictive planning enables preference alignment, the operational sequence of the framework, and how final translations are produced.
@@ 3.1 Test-Time Preference Alignment via Predictive Planning
Our key idea is to recast MT as trajectory optimization and thereby employ predictive planning for iterative translation refinement.
Machine Translation as Trajectory Optimization.
@@
One can interpret translation as an instance of trajectory optimization by (i) treating the source paragraph as the initial state $s_{1}$, (ii) treating each $a_{t}$ as the $t$-th segment of the translation, and (iii) setting $s_{t+1}$ as the concatenation of the source paragraph and the translations up to the $t$-th segment.
@@
Predictive Planning for Preference Alignment.
In paragraph-level MT, we adopt the principle of predictive planning in MPC for preference alignment. Figure 2 illustrates the framework of Plan2Align. Specifically: (i) We train a reward model on preference data. This reward model, which plays the role of utility function in MPC, assesses translation context quality and serves as a guiding function to align paragraph-level translations with preferences. By integrating MPC, we iteratively refine translations, ensuring alignment with both linguistic quality and contextual consistency. (ii) The LLM itself can serve as a predictive dynamics model as the next state can be obtained by appending the newly generated translation segment to the current state. (iii) To enable MPC for paragraph-level MT while ensuring strong context-level performance, we further introduce two key mechanisms for accumulating and utilizing high-quality translation experiences as described in the next two subsections.
3.1.1 Context Alignment and Context Buffer
A fundamental challenge in using MPC in paragraph-level MT is: How to continuously accumulate and retain valuable translation experience? A high-scoring translated paragraph does not necessarily indicate that every part of it is of high quality. Moreover, non-extensive LLMs struggle to identify and leverage high-quality contexts from translated paragraphs due to their limited long-text comprehension capabilities.
Instead of direct MPC application in MT, we modify MPC to selectively retain high-quality contexts from multiple paragraph-level translations instead of relying on a single best trajectory. This ensures effective accumulation and utilization of translation experience. To achieve this, we establish a context buffer, which stores high-quality context extracted from multiple paragraph-level translation experiences.

### v2

# Plan2Align: Predictive Planning Based Test-Time Preference Alignment for Large Language Models

@@ Abstract
Aligning Large Language Models with Preference Fine-Tuning is often resource-intensive. Test-time alignment techniques that do not modify the underlying model—such as prompting and guided decoding—offer a lightweight alternative. However, existing test-time alignment methods primarily improve short responses and fail to ensure coherence over extended contexts due to the myopic nature of token-level alignment. Moreover, these methods often incur a slowdown during inference. To address these challenges, we propose Plan2Align, a test-time alignment framework that formulates text generation as a predictive planning problem. Plan2Align adapts Model Predictive Control (MPC) to iteratively refine output by rolling out multiple complete responses and optimizing each segment. To more rigorously evaluate the effectiveness and efficiency, we focus on the more challenging task of long-text generation. Experiments on the long-form response subset of the HH-RLHF dataset and the WMT’24 Discourse-Level Literary Translation demonstrate that Plan2Align significantly enhances the performance of base LLMs. Compared to existing training-time and test-time alignment methods on LLaMA-3.1 8B, Plan2Align achieves comparable or superior results, while also delivering improved inference efficiency relative to prior test-time alignment approaches.
footnotetext: Equal contribution. Correspondence to Kuang-Da Wang <gdwang.cs10@nycu.edu.tw>, Ping-Chun Hsieh <pinghsieh@nycu.edu.tw>, and Chao-Han Huck Yang <hucky@nvidia.com>.
@@ 1 Introduction
Figure 1: Plan2Align mirrors control rollouts, selecting optimal segments for alignment.
The emergence of Large Language Models (LLMs), such as the GPT series (Brown et al., 2020; Achiam et al., 2023), LLaMAs (Touvron et al., 2023a, b), and Gemma (Team et al., 2024), has demonstrated remarkable efficacy in a wide range of NLP tasks (Stiennon et al., 2020; Hendrycks et al., 2021; Srivastava et al., 2023; Yu et al., 2024; Zhong et al., 2024). While these models exhibit strong performance out of the box, aligning their outputs to human preferences remains critical, especially for smaller-scale LLMs. For instance, in machine translation (Alves et al., 2024), smaller LLMs (e.g., under 10B parameters) frequently suffer from hallucinations, omissions, and semantic drift (Wu et al., 2024). Thus, aligning LLM outputs to preferences remains an essential yet challenging problem. Methods such as Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022) and Direct Preference Optimization (DPO) (Rafailov et al., 2023) have shown promising results in aligning preferences. However, these methods are typically resource-intensive.
Consequently, there is growing interest in test-time alignment techniques, which operate without modifying model parameters through strategies like prompting (Lin et al., 2024) or guided decoding (Khanov et al., 2024; Wang et al., 2024b; Li et al., 2024; Kong et al., 2024). However, existing methods face two fundamental challenges:
@@
    Limited Adaptivity: Many methods rely on guided decoding or value estimation, predicting outputs without the ability to revise earlier errors. This makes them brittle when the generation drifts from preferred trajectories, as they lack mechanisms for correction.
@@
    High Inference Overhead: Due to their focus on token or short-span alignment, these methods frequently introduce significant inference overhead, limiting their efficiency.
To address these challenges, we propose Plan2Align, a novel test-time replanning framework that formulates preference alignment as a predictive planning problem. Inspired by the principle of Model Predictive Control (MPC), a classic predictive planning method for trajectory optimization (Rao, 2014) widely used in robotics (Tassa et al., 2012, 2014; Nagabandi et al., 2020) and motion planning (Ji et al., 2016). As a predictive planning method, MPC optimizes a sequence of actions by continuously updating future plans based on past observations (Camacho and Bordons, 2007; Kouvaritakis and Cannon, 2016). Analogously, Plan2Align treats text generation as a trajectory in which each segment of output informs the next through an iterative feedback process. This framing enables the model to revise responses at the segment level [^1] The definition of a segment may vary depending on the task. For instance, in long-text translation, a segment may consist of three sentences, whereas in HH-RLHF, segments are defined by response position: the first third, middle third, and final third. The core idea is to leverage informative segments to guide the generation of more coherent and aligned outputs., improving both consistency and efficiency over long-form outputs. Figure 1 illustrates this correspondence, highlighting the intuition behind our approach.
Specifically, we introduce two core mechanisms:
  1. 1.
    Model-Predictive Segment Selection: We extend the MPC principle to selectively retain high-quality segments from multiple full responses using a reference buffer, rather than relying on a single best trajectory. This enables more effective accumulation and utilization of alignment-relevant signals over the course of generation.
  2. 2.
    Rewrite-Guided Replanning: To better guide small-scale LLMs toward aligned responses, we provide richer signals by recasting the task as a rewriting task. This structured prompting mechanism enables each iteration to build upon prior high-quality segments—not limited to the initial prompt but drawn from different locations in the response. By grounding future generation in selected high-quality text spans, our approach reduces reasoning difficulty for the model and promotes output consistency.
In this manner, Plan2Align enables efficient alignment and content refinement over a limited number of iterations, offering a practical and effective solution to preference alignment in small-scale LLMs.
Our contributions are summarized as follows:
  • We introduce a novel test-time alignment framework, transforming the alignment task into a predictive planning problem leveraging MPC principles, thus ensuring effective alignment.
  • We propose Plan2Align, integrating Model-Predictive Segment Selection and Iterative Replanning Generation, enabling efficient incremental improvements of generated text through effective utilization of high-quality segments within limited iterations.
  • We empirically evaluate Plan2Align using LLaMA-3.1 8B on challenging datasets including the WMT’24 discourse-level literary translation Benchmark and the long-form response subset of HH-RLHF. Our method demonstrates substantial improvements, achieving results comparable or superior to training-time alignment methods and outperforming existing test-time alignment approaches.
@@ 2 Related Work
2.1 Preference Alignment through Fine-tuning
Aligning large language models (LLMs) with human preferences has traditionally relied on post-training strategies. Supervised fine-tuning (SFT) (Ziegler et al., 2019) and reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) are widely used but computationally expensive. Direct Preference Optimization (DPO) (Rafailov et al., 2023) simplifies RLHF by converting it into a supervised learning objective, though it requires managing dual policies. More recent approaches like SimPO (Meng et al., 2024) and Contrastive Preference Optimization (CPO) (Xu et al., 2024) reduce memory and resource demands using reference models and contrastive signals. Despite these improvements, fine-tuning methods remain rigid and slow to adapt to changing data or objectives, posing challenges in dynamic environments.
2.2 Test-Time Preference Alignment
Test-time alignment offers a lightweight and adaptable alternative to fine-tuning by guiding generation during inference without updating model parameters. Beyond prompting and in-context learning, guided decoding is a major category of test-time alignment methods that explicitly control the generation process using external signals. ARGS (Khanov et al., 2024) is a representative example that incorporates reward model guidance at the token level, and InferAligner (Wang et al., 2024b) adopts a similar strategy. Among guided decoding methods, there are also approaches that directly modify internal representations. For instance, RE-Control (Kong et al., 2024) trains a value function on hidden states using the Bellman equation, and applies gradient-based optimization to align with preferences. TreeBoN (Qiu et al., 2024) and RAIN (Li et al., 2024) leverage tree-based structures: TreeBoN combines tree search with Best-of-N sampling, while RAIN performs self-evaluation without relying on a reward model to align preferences.
While these methods enhance alignment, they often incur high latency and computational overhead due to token-level control, especially in long-form generation. Moreover, value-guided decoding approaches struggle with inaccurate value estimation and instability when generation drifts from preferred trajectories. We propose a test-time alignment method that balances efficiency and alignment quality, achieving robust performance even in long-form generation settings.
@@ 3.1 Test-Time Preference Alignment via Predictive Planning
Our key idea is to recast preference alignment as trajectory optimization and thereby employ predictive planning for iterative refinement.
Text Generation Optimization as Trajectory Optimization.
@@
One can interpret text generation as an instance of trajectory optimization by (i) treating the input prompt as the initial state $s_{1}$, (ii) treating each $a_{t}$ as the $t$-th segment of the response, and (iii) setting $s_{t+1}$ as the concatenation of the input prompt and the response up to the $t$-th segment.
@@
Model Predictive Path Integral (MPPI).
MPPI (Williams et al., 2015) is a stochastic, sampling-based instantiation of model predictive control that reformulates the optimization in Eq. (2) as a path integral over trajectories. Rather than relying on gradient-based updates, MPPI approximates the control distribution by generating multiple candidate trajectories with random perturbations and computing their expected utility. Given a predictive model of environment dynamics, MPPI samples $K$ control sequences $\{(\boldsymbol{a}_{t:t+H}^{(i)}\}_{i=1}^{K}$ from a proposal distribution (typically Gaussian) and simulates the corresponding state trajectories. Each trajectory is scored by a cost or reward function $\mathcal{J}$, and the final control is computed as a weighted average:
$\boldsymbol{a}_{t}=\frac{\sum_{i=1}^{K}\exp\left(-\frac{1}{\lambda}\mathcal{J}^{(i)}\right)\boldsymbol{a}_{t}^{(i)}}{\sum_{i=1}^{K}\exp\left(-\frac{1}{\lambda}\mathcal{J}^{(i)}\right)},$ (3)
where $\lambda$ is a temperature parameter controlling the trade-off between exploration and exploitation. Unlike deterministic MPC, MPPI enables smoother updates by aggregating over multiple high-reward rollouts rather than relying on a single best trajectory.
Figure 2: Overview of Plan2Align. Left: Plan2Align adapts MPPI control to language generation by evaluating response segments with a reward model and reusing high-quality ones from a reference buffer. Right: We implement this with Model-Predictive Segment Selection (purple), where segment-level rollouts are scored and stored in a buffer, and Rewrite-Guided Replanning (gray), which uses aligned high-quality segments as rewriting task to progressively improve output.
Adapting MPPI for Preference Alignment.
Plan2Align draws inspiration from MPPI control and reinterprets its core principles for test-time generation with LLMs. As shown in Figure 2, we align components of MPPI with analogous operations in LLM-based generation:
  • In MPPI, a policy and world model are used to sample multiple future trajectories, each representing a sequence of actions. In Plan2Align, the LLM plays both roles, generating diverse candidate responses conditioned on an input prompt. These responses serve as discrete trajectories in language space.
  • Instead of evaluating entire trajectories, Plan2Align decomposes responses into segments and performs fine-grained segment-level evaluation using a reward model trained on preference data. This segment-wise assessment replaces MPPI’s scalar cost for each rollout, allowing more localized credit assignment and reuse of high-quality spans.
  • Rather than selecting one best trajectory or optimizing a control policy, Plan2Align stores high-scoring segments into a reference buffer. This buffer is used to construct rewriting prompts, enabling the LLM to adjust its behavior over successive iterations without modifying model parameters.
This adaptation retains the predictive planning spirit of MPPI—roll out, evaluate, and replan—but translates it into a discrete, segment-level alignment framework suited for long-form generation. As shown in the right half of Figure 2, rewriting prompts are constructed from a reference buffer of high-quality segments, allowing the model to iteratively improve its outputs during inference. The mechanisms that support this loop are described in the next two subsections.
3.2 Model-Predictive Segment Selection for Managing High-Quality Experience
We adapt the MPPI paradigm by selectively retaining high-quality segments from multiple responses, regardless of their position, instead of relying solely on the leading segment of the best-scoring output. This approach enables more efficient accumulation of useful generation trajectories. We implement this using a reference buffer that stores top-scoring segments across responses. Given that segment positions are fixed within a response, we maintain a buffer of size $k$ for each position, preserving the top-$k$ segments to guide subsequent replanning steps.
3.3 Rewrite-Guided Replanning for Improving Outputs
While high-quality experience can be retained through the reference buffer, effectively leveraging it to guide LLMs toward preference-aligned outputs remains a non-trivial challenge. This is particularly true for small-scale LLMs, which have limited capacity for contextual reasoning. Reducing the complexity of the replanning process is therefore critical.
To address this, we reframe the replanning task as a rewriting task. By providing LLMs with pre-aligned segments as input, the model can refine responses using high-quality content from previous iterations. Before rewriting, the system filters and retains only those segments that exceed a predefined quality threshold [^2] The threshold is typically based on the reward score of a preferred response. If the threshold is set too high, buffer updates may occur too slowly. If set too low, the buffer may include suboptimal segments.. This ensures that only informative and preference-aligned segments are carried forward, preventing contamination from low-quality content.
In this way, the model is not required to infer the exact insertion points for each segment. Instead, segments are positioned at fixed locations, and their combination serves as the basis for rewriting. This design lowers the reasoning burden while still allowing the model to produce improved responses. These refined responses are then used to update the buffer in the next iteration, promoting stable improvements in generation quality and alignment with preferences.
3.3.1 Accelerating the Preference Alignment Process
To accelerate preference alignment, Plan2Align aims to minimize the number of rewriting iterations by efficiently discovering diverse high-quality segments. We introduce stochasticity into rewriting task construction to promote exploration and increase the likelihood of retrieving useful content. Below, we describe how segments are defined and sampled for two different tasks. The corresponding rewriting prompt templates for each task are provided in Appendix G.1.2 and Appendix G.2.2.
Response Generation.
Response generation offers a relatively simple setting. The LLM generates $n$ responses per iteration, and each is divided into $m$ sentence-level segments of equal length. Segment positions are fixed to simplify alignment. Because the reward model assigns a single score per response, we uniformly propagate the score to all its segments. To construct rewriting prompts, we sample one segment per position from the top-$k$ highest scoring segments in the buffer. This randomness encourages diversity and helps discover better segments. If a sampled segment falls below a predefined quality threshold, we omit it, prompting the model to refine that portion.
Translation.
In translation, segment alignment is naturally guided by source–target correspondence. We define source-based segments as three-sentence spans with overlap via a sliding window (e.g., sentences 1–3, 2–4). Unlike response generation, the reward model can independently evaluate each segment, enabling more precise quality assessment and a reference buffer that better captures high-quality translations. Translation introduces additional challenges due to alignment noise between source and target segments, which can degrade prompt quality. To address this, we sample a subset of segments during intermediate iterations and progressively increase the sample size. The output translation uses all buffered segments. See Appendix Algorithm 1 for details.
4 Experiments
We evaluate Plan2Align on two challenging long-form generation tasks—long-form response generation and paragraph-level machine translation—to better reveal performance differences.
4.1 Preference Dataset
4.1.1 Long-Form Response Dataset
For long-form response generation, we use the Dahoas/full-hh-rlhf [^3] https://huggingface.co/datasets/Dahoas/full-hh-rlhf dataset, which is widely adopted for LLM alignment. This dataset is designed to improve AI assistant behavior in terms of helpfulness and harmlessness. Each sample consists of a prompt and two responses, with one labeled as preferred based on human judgments. Since the response lengths in the dataset vary significantly, we select samples based on the length of the chosen responses. Specifically, we construct the training set using the top 6K samples with the longest chosen responses from training set, and using the top 1024 longest chosen responses from the testing set to construct test set.
4.1.2 Paragraph-Level Translation Dataset
To establish a suitable preference dataset for long-text MT, we use the latest WMT’24 Discourse-Level Literary Translation benchmark (Wang et al., 2024a) for our experiments. This dataset allows for evaluating how well translation methods preserve contextual relationships during translation. The available language pairs include: Chinese →English, Chinese →German, and Chinese →Russian. To keep inputs within the context window of our chosen LLMs, we segment each training and validation instance into chunks of up to 1024 tokens, using GPT-4’s tokenizer to standardize token length. This preprocessing ensures that paragraph-level MT tasks are handled within the model’s constraints.
Preference Data Generation.
The preference data is derived from the training set of the dataset. Specifically, each instance is segmented into paragraphs of up to 1024 tokens. From each translation direction, we sample 2K paragraphs, resulting in a total of 6K paragraphs for constructing the preference data. We generate translation outputs using LLaMA-3.1-8B-Instruct, Gemma-2-9B, and GPT-4o. The translations are then evaluated using MetricX-24-XL (Juraska et al., 2024), a reference-free quality estimation (QE) model. The QE scores help us select high-quality translations: the translation with the highest score is designated as the preferred translation, while the one with the lowest score is considered not preferred. The translation with a middle score is disregarded.
4.2 Reward Model Training
Preference data is crucial for training alignment methods and for training the reward model used in Plan2Align. The collected preference data is utilized to train the reward model in Plan2Align, allowing it to evaluate the quality of contextual responses. The reward model trained for long-form responses achieves a validation accuracy of 83.78%, while the model trained for paragraph-level translation attains a validation accuracy of 88.53%. Further details on the formation of preference data can be found in Appendix C, and detail of training can be found in Appendix F.2.
4.3 Evaluation Metrics
4.3.1 Long-Form Responses
Drawing on prior work (Khanov et al., 2024; Kong et al., 2024), we evaluate responses quality using the following five metrics:
Average Reward reflects the average score assigned by the reward model during Plan2Align and ARGS inference, which indicates alignment with helpfulness and harmlessness preferences. Coherence is measured as the cosine similarity between the embeddings of the prompt and its continuation, computed using a SimCSE model (Gao et al., 2021), capturing how well the generated response maintains topical consistency. Diversity assesses lexical variation in the output, calculated by aggregating n-gram repetition rates. For a given continuation $y$, the diversity score is defined as $\text{diversity}(y)=\prod_{n=2}^{4}\frac{\text{unique $n$-grams}(y)}{\text{total $n$-grams}(y)}$, with higher values indicating broader vocabulary use. Win Rate captures the proportion of pairwise comparisons in which a model’s response is rated higher than a reference response by GPT-4 (OpenAI et al., 2024). Following the ARGS evaluation protocol (Khanov et al., 2024), GPT-4 is prompted to evaluate overall response quality, considering helpfulness, harmlessness, relevance, accuracy, depth, creativity, and detail. The full evaluation prompt is included in Appendix E. Lastly, Inference Time refers to the average time (in seconds) required to generate a single response, reflecting the computational efficiency of each method.
4.3.2 Paragraph-Level MT
In machine translation tasks, most evaluations rely on existing model-based metrics. Following CPO (Xu et al., 2024), we adopt COMET [^4] Unbabel/wmt22-comet-da and COMET-KIWI [^5] Unbabel/wmt22-cometkiwi-da (the reference-free version) for our evaluation. While model-based metrics have taken over string-based metrics as the state-of-the-art, all model-based metrics (e.g. COMET) are still only trained with sentence-level data, and machine translation evaluation on paragraph-level remains an open question. A prior work Deutsch et al. (2023) argued that existing metrics generalize well to paragraph-level, but the test set in that work is a mere concatenation of existing datasets translated by sentence-level models, and does not take into account the context-level improvements we are aiming at. In fact, in our preliminary study, we observed serious pathological behaviors of these metrics when evaluating our paragraph-level outputs (see Appendix H.2). This calls for a significant rework of the existing evaluation paradigm. The changes we’ve made to our machine translation evaluation paradigm are as follows:
Contextual Sliding Window.
Instead of simply feeding the source, translation and reference sentences into COMET and COMET-KIWI, we follow Vernikos et al. (2022) and add aligned and concatenated context as inputs to the model-based metrics. We call these metrics CW-COMET and CW-KIWI. We use the same size of 3 as described in Section 3.3.1.
Sentence Segmentation and Alignment.
Similar to IWSLT evaluation paradigm for speech translation Ahmad et al. (2024), we propose to treat sentence segmentation and alignment as part of the evaluation for our paragraph-level evaluation. We follow the same general procedure as described in Section 3.3.1. We handle sentence alignment differently according to whether the metric requires reference input. Misaligned sentences remain in their original positions, with empty strings as placeholders where necessary. If a sliding window contains only empty strings on the one side, a default score of zero is given. This penalizes over-translation and omission, aligning with our research objectives. We verified the validity of our paradigm via a sanity check. Details are in Appendix H.
4.4 Baselines
To ensures consistency in comparing the effectiveness, we evaluate all training-time alignment methods on LLaMA-3.1-8B-Instruct and also use it as the backbone model for test-time alignment methods, including Plan2Align. Additionally, we compare these results against the original LLaMA-3.1-8B-Instruct to further assess improvements and changes, serving as the baseline for performance. Implementation details of Plan2Align, including parameters and prompt design, are in Appendix G.
Test-Time Alignment Models.
We select ARGS (Khanov et al., 2024) and RAIN (Li et al., 2024) as they both represent test-time alignment methods similar to Plan2Align but with distinct operational mechanisms. To better isolate the effect of predictive planning, we also compare Plan2Align with Best-of-N sampling. Additionally, we introduce a simplified variant of our method, referred to as Vanilla MPC. This ablation removes both the segment-level evaluation and the use of a reference buffer. Instead of constructing rewriting tasks, Vanilla MPC simply selects the best-scoring response or translation from each iteration and feeds it into the next generation round for refinement. Implementation details of all test-time alignment baselines are provided in Appendix F.1.
Training-Time Alignment Models.
We also compare Plan2Align with training-time alignment methods. We consider SFT directly on the same preference dataset, which is typically a strong baseline for translation tasks. Moreover, we include SimPO for comparison as it serves as a strong benchmark for training-time alignment in general. We also consider DPO as it serves as a mainstream representative of the training-time-alignment series, with training details available in Appendix F.2.
GPT-4o and Qwen-2.5-14B.
For paragraph-level MT, we select GPT-4o as the upper bound for performance. GPT-4o is not optimized for translation, but closely approaches that of specialized translation systems (Shahriar et al., 2024). We also include Qwen-2.5-14B, as it has demonstrated strong performance in Chinese language tasks and serves as a representative open-source alternative.